OpenStudy (anonymous):

csc(cos^-1 7/25) is the answer 25/24??

OpenStudy (anonymous):

How can you get this without a calculator?

OpenStudy (anonymous):

Assume cos^-1 7/25 = theta

OpenStudy (anonymous):

this says basically if cosine of your number is 7/25, then what is the cosecant of your number cosine is 7/25, by pythagoras sine is 24/25 so yes, that means cosecant is 25/24

OpenStudy (anonymous):

cos theta = 7/25

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what's csc theta?

OpenStudy (anonymous):



OpenStudy (anonymous):

ohh I need to put cos 7/25 in pyth identity

OpenStudy (anonymous):

get sin, then put 1 over sin

OpenStudy (anonymous):

yes, then find sin theta and take the reciprocal

OpenStudy (anonymous):

right, or even hypotenuse over opposite if you draw a triangle

OpenStudy (anonymous):

omg! so easy. Thanks guys!

OpenStudy (anonymous):

i wish my math book said that

OpenStudy (anonymous):

only difference between these problems and ones you have done before is the wording. now you are asked for \(\csc(\cos^{-1}(x))\) before you were asked "if cosine x = something, then what are the other trig functions?" it is the same problem really

OpenStudy (anonymous):

that makes total sense, thanks